# Steering Embedding Models with Geometric Rotation: Mapping Semantic Relationships Across Languages and Models

### v1

@@ D.1 Diversity Controls
D.2 Feature-based Transformation Methodology
@@ D.2.1 Transformation Procedure
D.3 Feature-Controlled Prompting
To generate each transformation in a systematic and reproducible manner, we employ a feature-controlled prompting strategy with a large language model (LLM). Each prompt is carefully templated to specify the source language, the desired transformation type, and a set of fine-grained feature tags that guide the model’s output. For example, a prompt might indicate the language code (”[TA]” for Tamil), the transformation (”Politeness Rephrase”), and a particular strategy or keyword (such as ”add honorific”) relevant to that transformation. By explicitly encoding these features, we ensure that the LLM produces the intended variation—whether a more polite rephrasing, a negated statement, or a conditional construction—in a consistent and transparent way.
To further guarantee balanced coverage, we maintain a metadata table that enumerates all possible sub-types or strategies for each transformation. This enables us to stratify the sampling of transformation features across languages and sentences, ensuring that every variant type is equally represented. For instance, multiple politeness strategies (e.g., adding honorifics, using indirect language) or different negation words (”no” vs. ”not”) are distributed uniformly across the dataset. This controlled coverage is critical for fair comparisons: it prevents any language from being overrepresented by a particular style of rephrasing or negation, and minimizes inadvertent correlations between language and transformation realization. Our stratified sampling approach follows established principles of controlled experimental design, providing a robust foundation for cross-lingual embedding analysis.
All transformed sentences are generated using a single, consistent LLM—specifically, GPT-4.5—with a temperature of 1.0 and a maximum token limit of 128 per prompt. The relatively high temperature encourages diversity in phrasing, while the one-shot generation policy (taking the first model output without retries or manual curation) avoids selection bias. With carefully constructed prompts, the model reliably produces valid transformations on the first attempt, and all outputs remain in the target language specified by the prompt. This procedure ensures that our dataset is both systematically varied and reproducible, supporting rigorous downstream analysis.
D.4 Quality Control and Deduplication
@@
Beyond deduplication, we applied a suite of diversity controls to guard against artefacts arising from narrow lexical or topical coverage. Each neutral sentence prompt was drawn from a wide range of context categories and exemplars, ensuring topical breadth before any transformation was applied. Within each language, we shuffled sentence–feature assignments so that no specific lexical field correlated with a particular transformation subtype. For each transformation, we uniformly sampled property values (such as negation particles or politeness strategies) per language and sentence, guaranteeing that every combination of language and subtype appeared the same number of times. After generation, we removed near-duplicates and enforced a 5–25 token length window, which empirically yielded a near-uniform length distribution. Together, these steps ensure that our corpus varies in topic, syntax, and lexical choice while remaining balanced across languages and transformation subtypes, providing a robust foundation for subsequent embedding analysis.
D.5 Embedding Generation
@@
Our approach assumes that all sentence embeddings reside in a shared semantic space where linear structure is meaningful. We adopt the perspective that this space forms a latent manifold encoding universal semantic features, as hypothesized by Jha et al. (2025). In this framework, certain directions in the embedding space correspond to specific attributes, such as politeness or negation. If sentence transformations truly correspond to adding or subtracting a semantic attribute, we expect the difference vector (variant minus source) to be relatively consistent across examples. This aligns with the ”universal geometry for embeddings” framework, in which multilingual embeddings from different models or languages can be brought to a common representation where semantic differences are captured by geometric translations. While our work stays within a single encoder’s space, we leverage a similar idea: analyzing whether the transformation ”rotors” (difference vectors) cluster for similar transformations across languages. This methodology sets the stage for validating whether these quasi-linear transformations indeed behave like translations in a Riemannian semantic space (Jha et al., 2025), which we explore in the next section via rotor-based analysis of the embedding differences.
@@
D.6 Final Dataset Statistics
The resulting corpus comprises 1,000 neutral sentences in each of the seven languages, totaling 7,000 examples. For English neutral sentences, the mean token length is 9.1 tokens (with a median of 9.0 tokens), with token counts ranging from 3 to 12 tokens and an average character length of 54.4 characters. This distribution confirms that our generation process produced concise, natural sentences suitable for semantic transformation analysis across languages and transformation types.
@@ D.6 Final Dataset Statistics
Appendix E LLM Usage Disclosure
@@
Appendix F RISE vs Baseline Comparisons
This section presents comprehensive comparisons between RISE and random baseline prototypes to validate that RISE learns meaningful semantic directions rather than benefiting from arbitrary vector orientations.
F.1 Cross-Language Transfer vs Random Baselines
The following figures show detailed heatmaps comparing RISE performance against random prototypes of equivalent magnitude across all language pairs and phenomena. Each comparison uses 10,000 random trials to ensure statistical robustness.
F.2 RISE Baseline Comparison
Figure 7 demonstrates the baseline validity of RISE by comparing it against random prototypes across multiple language transfer scenarios. The consistent and substantial advantages (ranging from 5.1× to 26.2×) across all models and phenomena provide crucial validation that RISE learns meaningful semantic directions rather than exploiting statistical artifacts. Notably, cross-language transfers often maintain or even exceed monolingual performance relative to random baselines, confirming that RISE captures universal semantic patterns that generalize across language boundaries. Overall, RISE analyses show that embedding models encode some transformations as universal operators, but others remain highly culture- and resource-dependent. Future research should refine evaluation benchmarks to account for phenomenon-specific variability and investigate training regimes that promote balanced universality without sacrificing discriminative capacity.
Figure 7: RISE vs Random Baseline Comparisons across Language Transfer Scenarios.
Top: English monolingual analysis showing RISE performance vs random prototypes for all three models and phenomena.
Middle: English prototype → Spanish target cross-language transfer demonstrating maintained advantages over random baselines.
Bottom: Japanese prototype → English target transfer confirming universal semantic patterns across diverse language pairs. All comparisons use 10,000 random trials for statistical robustness.
@@ F.2 RISE Baseline Comparison
F.3 Phenomenon-Specific Performance vs Random Baselines
@@
These baseline comparisons provide crucial validation that RISE’s strong performance stems from learning meaningful semantic transformations rather than exploiting statistical artifacts or benefiting from arbitrary vector orientations in high-dimensional spaces.
F.4 Detailed Baseline Comparison Analysis
@@
Model Phenomenon RISE Perf Random Baseline Adv Ratio OpenAI (3072d) Conditionality 0.463 0.057 ± 0.0003 8.1× Negation 0.210 0.041 ± 0.0002 5.1× Politeness 0.181 0.031 ± 0.0002 5.8× BGE-M3 (1024d) Conditionality 0.610 0.057 ± 0.0003 10.7× Negation 0.391 0.041 ± 0.0002 9.5× Politeness 0.461 0.031 ± 0.0002 14.9× mBERT (768d) Conditionality 0.625 0.057 ± 0.0003 11.0× Negation 0.624 0.041 ± 0.0002 15.2× Politeness 0.294 0.031 ± 0.0002 9.5×
@@
Transfer Scenario OpenAI (3072d) BGE-M3 (1024d) mBERT (768d)

### v2

@@ E.1 Diversity Controls
E.2 Feature-based Transformation Methodology
@@ E.2.1 Transformation Procedure
E.3 Feature-Controlled Prompting
To generate each transformation in a systematic and reproducible manner, we employ a feature-controlled prompting strategy with a large language model (LLM). Each prompt is carefully templated to specify the source language, the desired transformation type, and a set of fine-grained feature tags that guide the model’s output. For example, a prompt might indicate the language code (“[TA]” for Tamil), the transformation (“Politeness Rephrase”), and a particular strategy or keyword (such as “add honorific”) relevant to that transformation. By explicitly encoding these features, we ensure that the LLM produces the intended variation—whether a more polite rephrasing, a negated statement, or a conditional construction—in a consistent and transparent way.
To further guarantee balanced coverage, we maintain a metadata table that enumerates all possible sub-types or strategies for each transformation. This enables us to stratify the sampling of transformation features across languages and sentences, ensuring that every variant type is equally represented. For instance, multiple politeness strategies (e.g., adding honorifics, using indirect language) or different negation words (“no” vs. “not”) are distributed uniformly across the dataset. This controlled coverage is critical for fair comparisons: it prevents any language from being overrepresented by a particular style of rephrasing or negation, and minimizes inadvertent correlations between language and transformation realization. Our stratified sampling approach follows established principles of controlled experimental design, providing a robust foundation for cross-lingual embedding analysis.
All transformed sentences are generated using a single, consistent LLM (GPT-4.5) with a temperature of 1.0 and a maximum token limit of 128 per prompt. The relatively high temperature encourages diversity in phrasing, while the one-shot generation policy (taking the first model output without retries or manual curation) avoids selection bias. With carefully constructed prompts, the model reliably produces valid transformations on the first attempt, and all outputs remain in the target language specified by the prompt. This procedure ensures that our dataset is both systematically varied and reproducible, supporting rigorous downstream analysis.
E.4 Quality Control and Deduplication
@@
Beyond deduplication, we applied a suite of diversity controls to guard against artifacts arising from narrow lexical or topical coverage. Each neutral sentence prompt was drawn from a wide range of context categories and exemplars, ensuring topical breadth before any transformation was applied. Within each language, we shuffled sentence–feature assignments so that no specific lexical field correlated with a particular transformation subtype. For each transformation, we uniformly sampled property values (such as negation particles or politeness strategies) per language and sentence, guaranteeing that every combination of language and subtype appeared the same number of times. After generation, we removed near-duplicates and enforced a 5–25 token length window, which empirically yielded a near-uniform length distribution. Together, these steps ensure that our corpus varies in topic, syntax, and lexical choice while remaining balanced across languages and transformation subtypes, providing a robust foundation for subsequent embedding analysis.
E.5 Embedding Generation
@@
Our approach assumes that all sentence embeddings reside in a shared semantic space where linear structure is meaningful. We adopt the perspective that this space forms a latent manifold encoding universal semantic features, as hypothesized by Jha et al. (2025). In this framework, certain directions in the embedding space correspond to specific attributes, such as politeness or negation. If sentence transformations truly correspond to adding or subtracting a semantic attribute, we expect the difference vector (variant minus source) to be relatively consistent across examples. This aligns with the “universal geometry for embeddings” framework, in which multilingual embeddings from different models or languages can be brought to a common representation where semantic differences are captured by geometric translations. While our work stays within a single encoder’s space, we leverage a similar idea: analyzing whether the transformation “rotors” (difference vectors) cluster for similar transformations across languages. This methodology sets the stage for validating whether these quasi-linear transformations indeed behave like translations in a Riemannian semantic space (Jha et al., 2025), which we explore in the next section via rotor-based analysis of the embedding differences.
@@
E.6 Final Dataset Statistics
The resulting corpus comprises 1,000 neutral sentences in each of the seven languages, totaling 7,000 examples. For English neutral sentences, the mean token length is 9.1 tokens (with a median of 9.0 tokens), with token counts ranging from three to 12 tokens and an average character length of 54.4 characters. This distribution confirms that our generation process produced concise, natural sentences suitable for semantic transformation analysis across languages and transformation types.
@@ E.6 Final Dataset Statistics
Appendix F LLM Usage Disclosure
@@
Appendix G Downstream Task Analysis
As requested by reviewers, we completed a downstream classification analysis. Due to time constraints, we focused on a single well-defined task: detecting negation in the English subset of the Synthetic Multilingual dataset. We evaluated how well a classifier trained on MDV-transformed and RISE-transformed sentences performed on a held-out test set of 1919 unpaired sentences (961 with negation, 958 without). The test set was generated with the same specifications described in Appendix D & E.
Now, both methods perform well on this task. MDV achieves strong recall (92.1%) and overall accuracy (87.2%), showing that even a simple mean displacement vector captures meaningful geometric regularities in the transformation. Yet, RISE yields a stronger downstream performance and outperforms MDV across all metrics (93.0% accuracy, 92.1% precision, 94.0% recall, and 93.0% F1). The positive results of both methods reinforces the broader claim that spherical, non-linear techniques are effective tools for capturing semantic-syntactic transformations in high-dimensional embedding spaces.
Method Accuracy Precision Recall F1
MDV 0.872 0.840 0.921 0.878
RISE 0.930 0.921 0.940 0.930
Table 9: Downstream negation classification performance for MDV and RISE transformations.
Appendix H RISE vs Random Baseline Comparisons
This section presents comprehensive comparisons between RISE and random baseline prototypes to validate that RISE learns meaningful semantic directions rather than benefiting from arbitrary vector orientations. The following figures show detailed graphs, heatmaps, and tables comparing RISE performance against random prototypes of equivalent magnitude across all language pairs and phenomena. Each comparison uses 10,000 random trials to ensure statistical robustness.
Figure 12 highlights three select language transfer scenarios and Figures 1315 demonstrate the baseline validity of RISE by comparing it against random prototypes across multiple language transfer scenarios. The consistent and substantial advantages (ranging from 5.1× to 26.2×) across all models and phenomena provide crucial validation that RISE learns meaningful semantic directions rather than exploiting statistical artifacts. Notably, cross-language transfers often maintain or even exceed monolingual performance relative to random baselines, confirming that RISE captures universal semantic patterns that generalize across language boundaries. Overall, RISE analyses show that embedding models encode some transformations as universal operators, but others remain highly culture- and resource-dependent. Future research should refine evaluation benchmarks to account for phenomenon-specific variability and investigate training regimes that promote balanced universality without sacrificing discriminative capacity.
Figure 12: RISE vs Random Baseline Comparisons across select language transfer scenarios.
Top: English monolingual analysis.
Middle: English prototype → Spanish target cross-language transfer.
Bottom: Japanese prototype → English target cross-language transfer.
@@ Appendix H RISE vs Random Baseline Comparisons
H.1 Phenomenon-Specific Performance vs Random Baselines
Figures 16 provide crucial validation that RISE’s strong performance stems from learning meaningful semantic transformations rather than exploiting statistical artifacts or benefiting from arbitrary vector orientations in high-dimensional spaces.
@@ H.1 Phenomenon-Specific Performance vs Random Baselines
H.2 Detailed Baseline Comparison Analysis
@@
Model Phenomenon RISE Perf Random Baseline Adv Ratio TE3L (3072d) Conditionality 0.463 0.057 ± 0.0003 8.1× Negation 0.210 0.041 ± 0.0002 5.1× Politeness 0.181 0.031 ± 0.0002 5.8× bge-m3 (1024d) Conditionality 0.610 0.057 ± 0.0003 10.7× Negation 0.391 0.041 ± 0.0002 9.5× Politeness 0.461 0.031 ± 0.0002 14.9× mBERT (768d) Conditionality 0.625 0.057 ± 0.0003 11.0× Negation 0.624 0.041 ± 0.0002 15.2× Politeness 0.294 0.031 ± 0.0002 9.5×
@@
Transfer Scenario TE3L (3072d) bge-m3 (1024d) mBERT (768d)